Hu if a,b,c are positive real no and a+b+c=1 then a/b+c + b/c+a + c/b+a is

Grade:11

1 Answers

Deepak Kumar Shringi
6 years ago
To solve this type of questions use short trick here put a=b=c=1/3
and then 1/2+1/2+1/2 = 3/2 now check options